OpenStudy (anonymous):

Using the given zero, find one other zero of f(x). Explain the process you used to find your solution. 2 - 3i is a zero of f(x) = x4 - 4x3 + 14x2 - 4x + 13.

OpenStudy (jdoe0001):

do you know how to do long division?

OpenStudy (jdoe0001):

ahemm. I meant, long division of polynomials that is

OpenStudy (anonymous):

kind of.. not really sure how they even got the i in 2-3i though.

OpenStudy (jdoe0001):

well.. is a complex root, that is, an imaginary expression, and in this case, it's given

OpenStudy (anonymous):

oh okay. so how do i find the other zero?

OpenStudy (jdoe0001):

ok... well, the complex root comes from using the the square root from say the quadratic formula, like \(\bf x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\) notice the \(\boldsymbol{\pm}\) which means it has a "negative version" of the root, and a "positive version" of it so if we see a complex one, we can tell that it's not there all by her lonesome, it has a companion, her conjugate

OpenStudy (jdoe0001):

so... gimme a sec

OpenStudy (jdoe0001):

\(\bf 2 - 3i \quad conjugate \implies 2+3i\\ \textit{so we can say that our roots are}\\ x = 2 - 3i\implies x - 2 + 3i=0\implies (x - 2 + 3i)=0\\ \quad \\ x = 2 + 3i\implies x - 2 - 3i=0\implies (x - 2 - 3i)=0\\ \quad \\ \textit{thus}\\ \quad \\ (x - 2 + 3i)(x - 2 - 3i)=0\\ \quad \\ \textit{recall that }\quad (a-b)(a+b) = (a^2-b^2)\\ \quad \\ (x - 2 + 3i)(x - 2 - 3i)=0\implies (x - 2) + (3i)(x - 2) - (3i)=0\\ \quad \\ (x - 2)^2 - (3i)^2\implies x^2-4x+2^2-3^2i^2\implies x^2-4x+4-9(-1)\\ \quad \\ x^2-4x+9\implies x^2-4x+13\)

OpenStudy (jdoe0001):

so the quotient of dividing our polynomial, for THAT root, will give us a quadratic, that has our other 2 roots

OpenStudy (jdoe0001):

hmm... I'm not getting 0 as remainder

OpenStudy (debbieg):

Wait a sec.... if you have one complex root, then its complex conjugate has to be another root. So "find one other zero".... just give the complex conjugate as another zero.

OpenStudy (jdoe0001):

hmmm done that

OpenStudy (jdoe0001):

ohhh... I see shoot

OpenStudy (jdoe0001):

they don't want ALL ROOTS, just some OTHER ROOT

OpenStudy (debbieg):

lol yes, I see that... just wans't sure why we were worrying about doing long division of the polynomial. :)

OpenStudy (jdoe0001):

har har har.. well, yes I was =)

OpenStudy (debbieg):

Still always a valuable exercise. :)

OpenStudy (jdoe0001):

hmmm,, it has not real roots, no wonder the polynomial wasn't giving me a 0 remainder heheh

OpenStudy (jdoe0001):

well, that's what I get for not doing a rule of signs check first =)

OpenStudy (anonymous):

i don't get it. so x^2-4x+13 is the answer?

OpenStudy (jdoe0001):

hmmm no, see what they're asking is "Using the given zero, find \(\bf \text{one other zero}\) of f(x). Explain the process you used to find your solution."

OpenStudy (jdoe0001):

so, as I explained before, complex roots don't come all by their lonesome, they have their conjugate companion so the \(\bf \text{one other zero}\) will be its conjugate :)

OpenStudy (anonymous):

so 2 + 3i?

OpenStudy (jdoe0001):

yeap